### v3

# Multiscale Analysis of the Conductivity in the Lorentz Mirrors Model

Abstract
We study deterministic transport in a random medium using the mirrors model, a lattice Lorentz gas at unit density in which a particle moves deterministically in a frozen random configuration of mirror–scatterers. Despite the absence of chaos and the existence of infinitely many finite trapping loops, numerical evidence suggests that the model exhibits normal conductivity in $d=3$. We develop a multiscale expansion for the crossing probability of a slab of width $N$, showing that $C_{N}\sim\kappa/(\kappa+N)$ for large $N$ and computing the conductivity constant $\kappa$ through a renormalization procedure based on scale concatenation.
The key idea is that a slab of width $2^{n+1}$ may be decomposed into two independent slabs of width $2^{n}$, and that the crossing event can be expressed as a sum over trajectories that revisit the interface between the two halves. This gives rise to a recursion relation for the crossing probability. For the mirrors model, it involves a correction factor with respect to a reference Markovian process that encodes correlations between crossings at scale $2^{n}$. These correlations are controlled through a closure assumption whose structure is shaped by the hard–core exclusion inherent to the reversible deterministic dynamics. The main contribution comes from second–order returns, whose asymptotics we compute explicitly.
In $d=3$ the recursion takes the form
$\kappa_{n+1}=\kappa_{n}\Bigl(1+\alpha\,\frac{\kappa_{n}}{2^{n}}+o(2^{-n})\Bigr),\qquad\alpha\simeq 0.0374,$
leading to a finite limit $\kappa_{\infty}\simeq 1.5403$, in remarkable agreement with numerical simulations. This value is close to the conductivity of a non–backtracking random walk, suggesting that the large–scale behavior of the mirrors model is effectively Markovian even though the microscopic dynamics is fully deterministic.
1 Introduction
Transport in deterministic systems with quenched disorder plays a central role in statistical mechanics. A fundamental question is whether a macroscopic law such as Fick’s law can emerge from a microscopic dynamics that is neither stochastic nor chaotic. A paradigmatic example is the lattice Lorentz gas, where a particle moves deterministically in a fixed random configuration of scatterers. Even in this simple setting, understanding the emergence of diffusion and normal conductivity remains a major challenge.
The mirrors model, introduced in [8], provides a particularly striking instance of such a system. The scatterers are “mirrors” placed at the sites of $\mathbb{Z}^{d}$, each implementing a random reflection chosen uniformly among those satisfying reversibility and the no–U–turn constraint. The resulting dynamics is fully deterministic, time–reversible, and non–chaotic; in particular, infinitely many finite loops occur with positive probability. This leads to strong memory effects and non–Markovian behavior that severely complicate probabilistic analysis. In fact, even the question of whether trajectories are typically open or closed is unresolved in $d=3$ at full density.
Nevertheless, numerical experiments [3] suggest that a remarkable form of universality emerges at large scales: the model exhibits normal conductivity in $d=3$, with a crossing probability close to that of a non–backtracking random walk. This raises two natural questions: (i) how can a deterministic, non–chaotic system effectively behave as a Markov process at large scales? and (ii) can the conductivity be computed analytically?
The goal of this paper is to answer both questions through a multiscale analysis of the crossing probability of slabs. For a slab $\Lambda_{N}$ of width $N$, we denote by $C_{N}$ the probability (with respect to the disorder) that a particle entering on the left exits on the right. The conductivity at scale $N$ is defined as
$\kappa_{N}=\frac{NC_{N}}{1-C_{N}},$
and our aim is to understand the asymptotic behavior of $\kappa_{N}$ as $N\to\infty$.
Our analysis is based on a decomposition of a slab of width $2^{n+1}$ into two slabs of width $2^{n}$. For a non–backtracking random walk, independence yields an exact recursion for the crossing probability, from which one derives $\kappa_{N}\equiv d/(d-1)$. In the mirrors model, the left and right halves are independent, but a trajectory that repeatedly crosses the interface forces different crossing segments to satisfy deterministic compatibility constraints inside each half. These induced correlations between successive crossings form the main difficulty. A closure hypothesis is used to control these correlations. Under this decomposition, the correlation functions split naturally into two sectors: one in which correlations vanish identically due to hard–core constraints imposed by the reversible dynamics, and another in which one assumes increasing independence between trajectories as the scale grows.
We introduce quantities $\eta_{n}(l)$ that measure the deviation from independence for trajectories that revisit the interface l1l-1 times at scale $2^{n}$. The conductivity correction at scale $2^{n+1}$ is governed by a weighted sum of the $\eta_{n}(l)$, with weights decaying geometrically in $l$. The dominant contribution arises from $l=2$. We show that in $d=3$ this contribution converges rapidly to a constant that we compute numerically and bound analytically. The resulting recursion is
$\kappa_{n+1}=\kappa_{n}\Bigl(1+\frac{\kappa_{n}}{2^{n}}(\alpha+o(2^{-n}))\Bigr),\qquad\alpha\simeq 0.0374.$
This leads to a finite limit $\kappa_{\infty}\simeq 1.5403$, in excellent agreement with simulations and very close to the value $3/2$ for the non–backtracking random walk.
Outline of the paper. Section 2 defines the model and the crossing probability. Section 3 recalls the Markovian benchmark given by the non–backtracking random walk. Section 4 presents the multiscale recursion for the mirrors model. Section 5 contains the closure hypothesis and the derivation of the conductivity recursion. Section 6 analyzes the dominant second–order term and computes the limit conductivity. Appendix A contains the computation of $C_{1}$, and Appendices B and C summarize the correlation structure and bounds on the error terms.
Figure 1: Two-dimensional mirrors model on a portion of the lattice $\mathbb{Z}^{2}$. Mirrors (red) reflect the deterministic trajectory (blue). Periodic vertical boundary conditions are indicated by short arrows leaving and re-entering the domain.
2 Model and Definitions
We consider a subset of $\mathbb{Z}^{d}$ of the following form:
$\Lambda_{N}=\{1,\ldots,M\}^{d-1}\times\{1,\ldots,N\},$
with $M$ much larger than $N$, and with periodic or sufficiently large transverse extent so that boundary effects in directions orthogonal to $\mathbf{e}_{1}$ may be neglected.
Let $(\mathbf{e}_{1},\ldots,\mathbf{e}_{d})$ be the canonical basis of $\mathbb{R}^{d}$, and define the set of possible velocities
$\mathcal{P}=\{\pm\mathbf{e}_{1},\ldots,\pm\mathbf{e}_{d}\}.$
The motion of particles, travelling on the edges of $\mathbb{Z}^{d}$ with unit velocity, is described at every time $t\in\mathbb{N}$ by a point in the phase space
$\mathcal{M}_{N}=\{(\mathbf{q},\mathbf{p}):\mathbf{q}\in\Lambda_{N},\ \mathbf{p}\in\mathcal{P}\}.$
$\Lambda_{N}$$q_{1}=1$$q_{1}=N$$I_{N}^{+}$$I_{N}^{-}$
Incoming phase-space boundaries $I_{N}^{+}$ and $I_{N}^{-}$.
$\Lambda_{N}$$q_{1}=1$$q_{1}=N$$O_{N}^{-}$$O_{N}^{+}$
Outgoing phase-space boundaries $O_{N}^{-}$ and $O_{N}^{+}$.
Figure 2: The slab $\Lambda_{N}$ as a subset of $\mathbb{Z}^{d}$ (here represented in two dimensions), together with the incoming sets $I_{N}^{\pm}$ (left) and outgoing sets $O_{N}^{\pm}$ (right) in the direction $\mathbf{e}_{1}$. The underlying lattice structure is shown explicitly as the edges of $\mathbb{Z}^{2}$, and the thick dots mark the phase-space positions where the velocities are attached.
We define the outgoing boundaries corresponding to particles leaving the slab on the left or right with appropriate velocities:
$O^{-}_{N}$ $=\{(\mathbf{q}-\mathbf{e}_{1},-\mathbf{e}_{1}):q_{1}=1\},$ (1)
$O^{+}_{N}$ $=\{(\mathbf{q}+\mathbf{e}_{1},\mathbf{e}_{1}):q_{1}=N\},$ (2)
and $O_{N}=O_{N}^{+}\cup O_{N}^{-}$. Similarly, the incoming sets are
$I^{-}_{N}$ $=\{(\mathbf{q},-\mathbf{e}_{1}):q_{1}=N\},$ (3)
$I^{+}_{N}$ $=\{(\mathbf{q},\mathbf{e}_{1}):q_{1}=1\},$ (4)
with $I_{N}=I_{N}^{-}\cup I_{N}^{+}$.
On the vertices of $\Lambda_{N}$, we put “mirrors” that modify the orientation of the velocity of the particles. For each $\mathbf{q}\in\Lambda_{N}$, the action of a mirror on the velocity of an incoming particle is represented by a bijection $\pi(\mathbf{q};\cdot)$ of $\mathcal{P}$ into itself. It satisfies the reversibility condition:
$\pi(\mathbf{q};-\pi(\mathbf{q};\mathbf{p}))=-\mathbf{p},\qquad\forall\,(\mathbf{q},\mathbf{p})\in\mathcal{M}_{N},$
and the no U–turn property:
$\pi(\mathbf{q};\mathbf{p})\neq-\mathbf{p}.$
The dynamics is defined on $\mathcal{M}_{N}$ by
$F(\mathbf{q},\mathbf{p};\pi)=\bigl(\mathbf{q}+\pi(\mathbf{q};\mathbf{p}),\,\pi(\mathbf{q};\mathbf{p})\bigr).$ (5)
It is straightforward to check that the map $F$ is bijective between $\mathcal{M}_{N}$ and $(\mathcal{M}_{N}\backslash I_{N})\cup O_{N}$.
At each site of $\mathbb{Z}^{d}$, we pick independently a bijection $\pi(\mathbf{q};\cdot)$ with a uniform law over the set of bijections satisfying the reversibility and no U–turn constraints. For $d\geq 2$, there are (2d1)!!(2d-1)!! such bijections, which is also the number of pairings of $2d$ edges. We denote by $\mathbb{M}$ the corresponding probability law over the mirrors $\{\pi(\mathbf{q},\cdot):q\in\Lambda_{N}\}$, and by $\langle\cdot\rangle$ the average with respect to this law.
We define
[…]
Since $F$ is a bijection, for each $x$, there is exactly one $x^{\prime}$ such that $T(x,x^{\prime};\pi)=1$, and we have
$\sum_{x^{\prime}\in\mathcal{M}_{N}\cup O_{N}}T(x,x^{\prime};\pi)=1,\quad\forall x,\ \forall\pi.$ (7)
The slab $\Lambda_{N}$ itself may be viewed as a single deterministic scatterer: for a fixed configuration of mirrors, each incoming point $x\in I_{N}$ is mapped to a unique outgoing point in $O_{N}$. For any $x\in I_{N}$, define the exit time
$n_{x}:=\inf\{n\geq 0:F^{n}(x;\pi)\in O_{N}\},$
and the induced map
$F_{N}(x;\pi)=F^{n_{x}}(x;\pi).$ (8)
Define
$T_{N}(x,x^{\prime};\pi)=\begin{cases}1&\text{if }F_{N}(x,\pi)=x^{\prime},\\ 0&\text{otherwise}.\end{cases}$ (9)
Since $F$ is a bijection between $\mathcal{M}_{N}$ and $(\mathcal{M}_{N}\backslash I_{N})\cup O_{N}$ and $\Lambda_{N}$ is finite, for any $x\in I_{N}$, $n_{x}$ is finite and not larger than $2d|\Lambda_{N}|$. There can be no loop containing a point in $I_{N}$ because no point $x\in\mathcal{M}_{N}$ is mapped to $I_{N}$.
The single-point crossing probability is
$p_{N}(x,x^{\prime}):=\left\langle T_{N}(x,x^{\prime};\pi)\right\rangle.$ (10)
Our main object of study is the probability that a particle entering the volume $\Lambda_{N}$ on the left side at point ${\bf 1}:=(1,\ldots,1)$ crosses the volume and exits at some site $(N+1,q_{2},\ldots,q_{d})$:
$C_{N}:=\sum_{x^{\prime}\in O^{+}_{N}}\left\langle T_{N}(x,x^{\prime};\pi)\right\rangle,$
with $x=(\mathbf{1},\mathbf{e}_{1})$. We also introduce the conductivity at scale $N$:
$\kappa_{N}=\frac{NC_{N}}{1-C_{N}}.$
We will argue that $\kappa_{N}\to\bar{\kappa}$ as $N\to\infty$ with $0<\bar{\kappa}<\infty$.
3 Markovian Benchmark: Non–Backtracking Random Walk
We now consider the average of the transition rule with respect to the disorder:
$p(x,x^{\prime})=\langle T(x,x^{\prime};\pi)\rangle=\frac{1}{2d-1}\delta_{\mathbf{q}^{\prime},\mathbf{q}+\mathbf{p}^{\prime}}(1-\delta_{\mathbf{p},-\mathbf{p}^{\prime}}),$ (11)
which is a transition probability for a Markov chain $(X_{n})_{n\in\mathbb{N}}$ on $\mathcal{M}_{N}$. This is the transition kernel of a kinematic non–backtracking random walk.
Let $\widehat{C}_{N}$ be the corresponding crossing probability:
$\widehat{C}_{N}=\mathbb{P}[X_{t_{x}}\in O^{+}_{N}\mid X_{0}=x],$
for any $x\in I^{+}_{N}$, where $t_{x}=\inf\{n:X_{n}\in O_{N}\}$ is the exit time.
It is possible to compute $\widehat{C}_{N}$ explicitly and show
$\widehat{C}_{N}=\frac{\widehat{\kappa}_{0}}{\widehat{\kappa}_{0}+N},\qquad\widehat{\kappa}_{0}=\frac{d}{d-1}.$ (12)
To see this, define $\widehat{c}_{n}=\widehat{C}_{2^{n}}$ and
$\widehat{\kappa}_{n}=\frac{2^{n}\widehat{c}_{n}}{1-\widehat{c}_{n}}.$
A slab of length $N=2^{n+1}$ can be divided into two equal parts of length $2^{n}$, say $\Lambda_{1}$ and $\Lambda_{2}$. A path starting at $x\in I^{+}_{2^{n+1}}$ can exit in $O^{+}_{2^{n+1}}$ by crossing first $\Lambda_{1}$ and then $\Lambda_{2}$, or by revisiting the interface between them.
Independence yields
$\widehat{c}_{n+1}=\sum_{l=0}^{\infty}(\widehat{c}_{n})^{2}(1-\widehat{c}_{n})^{2l}=\frac{\widehat{c}_{n}}{2-\widehat{c}_{n}}.$ (13)
It is easy to check that $\widehat{\kappa}_{n+1}=\widehat{\kappa}_{n}$, and thus $\widehat{\kappa}_{n}=\widehat{\kappa}_{0}$ for all $n\geq 0$, and
$\widehat{c}_{n}=\frac{\widehat{\kappa}_{0}}{2^{n}+\widehat{\kappa}_{0}}.$
Computing $\widehat{c}_{0}$ gives
$\widehat{c}_{0}=\frac{1}{2d-1}+\frac{2d-2}{2d-1}\frac{1}{2}=\frac{d}{2d-1},$
and hence $\widehat{\kappa}_{0}=d/(d-1)$.
4 Multiscale Decomposition and Recursion
[…]
Figure 3: Multiscale composition in 3D: $2^{n}+2^{n}\!\to\!2^{n+1}$, then $2^{n+1}+2^{n+1}\!\to\!2^{n+2}$.
We now turn to the mirrors model, where the dynamics is non–Markovian. Define $c_{n}=C_{2^{n}}$ and
$\kappa_{n}=\frac{2^{n}c_{n}}{1-c_{n}}.$
As above, we split a slab of length $2^{n+1}$ into two equal parts of length $2^{n}$, denoted $\Lambda_{1}$ (left) and $\Lambda_{2}$ (right). For each part we define the incoming phase space volumes $\mathcal{I}^{\pm}_{i}$ and outgoing phase space volumes $\mathcal{O}^{\pm}_{i}$ in a way strictly analogous to the definitions in Section 2.
We let $t_{n}=T_{2^{n}}$ denote the indicator of crossing at scale $2^{n}$. A path crossing the full slab of length $2^{n+1}$ and revisiting the interface l1l-1 times can be written as
$t_{n+1}(x,x^{\prime},\pi)=\sum_{l\geq 1}\prod_{i=1}^{l}t_{n}(x_{i},y_{i};\pi_{1})\prod_{i=1}^{l}t_{n}(y_{i},x_{i+1};\pi_{2}),$
where $x_{1}=x$, $x_{l+1}=x^{\prime}$, and $\pi_{1},\pi_{2}$ are the restrictions of $\pi$ to $\Lambda_{1}$ and $\Lambda_{2}$. The sums over $x_{i}$ run over $x_{i}\in\mathcal{O}^{-}_{2}$, for $i=1,\ldots,l$, and those over $y_{i}$ run over $y_{i}\in\mathcal{O}^{+}_{1}$.
Using independence of $\pi_{1}$ and $\pi_{2}$ and taking expectation yields
$c_{n+1}=\sum_{l\geq 1}\sum_{x_{l+1}\in\mathcal{O}^{+}_{2}}\left\langle\prod_{i=1}^{l}t_{n}(x_{i},y_{i};\pi)\right\rangle\left\langle\prod_{i=1}^{l}t_{n}(y_{i},x_{i+1};\pi)\right\rangle,$ (14)
with $x_{1}=(\underline{1},\mathbf{e}_{1})$, and implicit summation over repeated intermediate indices.
We define the transition kernel at scale $n$:
$c_{n}(x,x^{\prime}):=\langle t_{n}(x,x^{\prime};\pi)\rangle.$ (15)
By translation invariance in the directions orthogonal to ${\mathbf{e}}_{1}$, the term corresponding to $l=1$ is $(c_{n})^{2}$.
We introduce
$\eta_{n}(l)=\frac{1}{c_{n}^{2}(1-c_{n})^{2(l-1)}}\sum_{x_{l+1}\in\mathcal{O}^{+}_{2}}\left\langle\prod_{i=1}^{l}t_{n}(x_{i},y_{i};\pi)\right\rangle\left\langle\prod_{i=1}^{l}t_{n}(y_{i},x_{i+1};\pi)\right\rangle.$ (16)
This measures how correlations at scale $2^{n}$ deviate from independent concatenation when a path revisits the interface l1l-1 times.
We write
$c_{n+1}=\sum_{l=1}^{\infty}c_{n}^{2}(1-c_{n})^{2(l-1)}\eta_{n}(l).$
Defining
$S_{n}=\sum_{l=1}^{\infty}(1-c_{n})^{2(l-1)}\eta_{n}(l),$
with $\eta_{n}(1)=1$, we have
$c_{n+1}=c_{n}^{2}S_{n}.$
We also observe
$S_{n}-(1-c_{n})^{2}S_{n}=1+\sum_{l=1}^{\infty}(1-c_{n})^{2l}(\eta_{n}(l+1)-\eta_{n}(l)),$
hence
$S_{n}=\frac{1}{1-(1-c_{n})^{2}}(1+\Delta_{n}),$
with
$\Delta_{n}=\sum_{l=1}^{\infty}(1-c_{n})^{2l}(\eta_{n}(l+1)-\eta_{n}(l)).$ (17)
The analog of (13) becomes
$c_{n+1}=\frac{c^{2}_{n}(1+\Delta_{n})}{1-(1-c_{n})^{2}}=\frac{c_{n}}{2-c_{n}}(1+\Delta_{n}).$
We are interested in the induced recursion for $\kappa_{n}$:
$\kappa_{n+1}=\frac{2^{n+1}c_{n+1}}{1-c_{n+1}}=\frac{2^{n}c_{n}(1+\Delta_{n})}{1-c_{n}-\frac{c_{n}}{2}\Delta_{n}}.$ (18)
5 Closure Assumption and Second-Order Terms
In the definition of $\eta_{n}(l)$ and hence $\Delta_{n}$, each term is the product of two factors, each representing the joint probability of $l$ given trajectories in the box of size $2^{n}$ entering and ending at fixed sites on the boundaries. In other words the crossing probability at scale $n+1$ depends on correlations between trajectories at scale $n$.
We introduce below a closure hypothesis on the joint probabilities. The basic idea is that trajectories that do not share an entry/exit point display increasing independence as the scale grows. The trajectories that do share entry/exit are either completely forbidden or infinitely forced to appear together. This allows the inductive computation of $\kappa_{n}$ at every scale. See appendix B for a full explicit description of the closure hypothesis.
In the expression for $\eta_{n}(l)$ we insert
$t_{n}(x,y;\pi)=c_{n}(x,y)+\delta t_{n}(x,y),$ (19)
for each factor in the products. This defines $\delta t_{n}$. The term with no $\delta t$ gives 1, so it is natural to define $\delta_{n}(l)=\eta_{n}(l)-1$.
Assuming absolute convergence of the series defining $\Delta_{n}$ and using $\eta_{n}(1)=1$, we find
$\Delta_{n}=\sum_{l=1}^{\infty}\left((1-c_{n})^{2l}-(1-c_{n})^{2(l+1)}\right)\delta_{n}(l+1)=c_{n}(2-c_{n})\sum_{l=1}^{\infty}(1-c_{n})^{2l}\delta_{n}(l+1).$ (20)
We focus on $\delta_{n}(2)$ after using (19). This term involves products of two expectations of products of two $t_{n}$’s divided by $c^{2}_{n}(1-c_{n})^{2}$. The first purely $c_{n}$ term is cancelled by the subtraction defining $\delta_{n}$. All terms containing a single $\delta t_{n}$ vanish because $\langle\delta t_{n}(x,y;\pi)\rangle=0$. Thus we are left with $\delta_{n}(2)=R_{1}+R_{2}$, where
$c_{n}^{2}(1-c_{n})^{2}R_{1}$ $=\sum_{x_{3}\in O_{2}^{+}}\langle\delta t_{n}(x_{1},y_{1})\delta t_{n}(x_{2},y_{2})\rangle\langle\delta t_{n}(y_{1},x_{2})\delta t_{n}(y_{2},x_{3})\rangle,$ (21)
$c_{n}^{2}(1-c_{n})^{2}R_{2}$ $=2\sum_{x_{3}\in O_{2}^{+}}\langle\delta t_{n}(x_{1},y_{1})\delta t_{n}(x_{2},y_{2})\rangle c_{n}(y_{1},x_{2})c_{n}(y_{2},x_{3}).$ (22)
Here $x_{1}=(\underline{1},\mathbf{e}_{1})$, and the sum over repeated indices is implicit.
We now state the closure assumption in the case $l=2$ and in the special case appearing in (21) and (22). For any $x_{1}\neq x_{2}$:
$\langle t_{n}(x_{1},y_{1})t_{n}(x_{2},y_{2})\rangle=(1-\delta_{y_{1}y_{2}})\,c_{n}(x_{1},y_{1})c_{n}(x_{2},y_{2})\bigl(1+O(h(n))\bigr),$ (23)
with $h(n)\to 0$ as $n\to\infty$. If $x_{1}=x_{2}$, then
$\langle t_{n}(x_{1},y_{1})t_{n}(x_{2},y_{2})\rangle=\delta_{y_{1}y_{2}}c_{n}(x_{1},y_{1}).$
The correlation functions represent correlations between two crossings in opposite directions. Using (23), one can show (see Appendix C) that
$|R_{2}|\leq\frac{O(h(n))}{2^{n}},$ (24)
consistent with numerical simulations, which indicate a decay of order $1/(2^{n})^{2}$.
We now turn to the more relevant term $R_{1}$. Writing $R_{1}=R_{11}+R_{12}$ and splitting the sum in (21) according to $y_{2}=y_{1}$ or $y_{2}\neq y_{1}$, the terms with $y_{2}\neq y_{1}$ can be shown (again using closure) to be of order $(h(n))^{2}$ and thus negligible. The dominant contribution is
$c_{n}^{2}(1-c_{n})^{2}R_{11}=\sum_{x_{3}\in O_{2}^{+}}c_{n}(x_{1},y_{1})\bigl(c_{n}(y_{1},x_{2})\bigr)^{2}c_{n}(y_{1},x_{3}),$
using invariance under time reversal of $c_{n}(\cdot,\cdot)$. By translation invariance in $y_{1}$, we have $\sum_{x_{3}\in O_{2}^{+}}c_{n}(y_{1},x_{3})=c_{n}$, and hence
$(1-c_{n})^{2}R_{11}=\sum_{x_{2}\in O_{2}^{-}}\bigl(c_{n}(y_{1},x_{2})\bigr)^{2}\leq\Bigl(\sum_{x_{2}\in O_{2}^{-}}c_{n}(y_{1},x_{2})\Bigr)^{2}=(1-c_{n})^{2}.$
Thus $R_{11}\leq 1$. Numerical computation shows rapid convergence of $(1-c_{n})^{2}R_{11}$ to a constant independent of $n$:
$(1-c_{n})^{2}R_{11}\to 0.018704\pm 10^{-5}.$
The contribution to $\Delta_{n}$ of the term $l=2$ reads
$\Delta_{n}(2)=c_{n}(2-c_{n})\bigl((1-c_{n})^{2}R_{11}+O(h(n))\bigr).$
$x_{1}$$y_{1}=y_{2}$$x_{2}$$y_{3}$
Figure 4: Trajectories contributing to $R_{1}1$ for $l=2$.
Using (18) and $c_{n}=\kappa_{n}/(2^{n}+\kappa_{n})$, we obtain
$\kappa_{n+1}=\kappa_{n}\left(1+\frac{\kappa_{n}}{2^{n}}(\alpha+o(2^{-n}))\right),\qquad\alpha=0.0374,$
where $\alpha=2(1-c_{n})^{2}R_{11}$. Iterating this relation (neglecting $o(2^{-n})$) from $\kappa_{n}=1.5397\pm 3\times 10^{-4}$ at $n=8$ yields
$\kappa_{\infty}=1.5403.$
The comparison between the measured ratio
$1+\Delta_{n}=\frac{c_{n+1}(2-c_{n})}{c_{n}}$
and the result of the above computation is given in Figure 5.
Figure 5: The measured ratio $1+\Delta_{n}=c_{n+1}(2-c_{n})/c_{n}$ is plotted in blue with a 95% confidence interval, while $1+c_{n}(2-c_{n})(1-c_{n})^{2}R_{11}$ is plotted in red. The contribution of $R_{2}$ lowers the red values significantly only for the two first values of $N$, the contribution of $R_{12}$ increases those values in a way that is almost not visible at this scale.
Figure 6: A plot of measured values of $\kappa$ as a function of the slab length displaying convergence to $\kappa_{\infty}$ computed above.
6 Discussion and Perspectives
The mirrors model is a deterministic dynamical system in a random environment, and as such it does not fall into the classical framework of stochastic homogenization. Nonetheless, the crossing probability $C_{N}$ exhibits the universal scaling law
$C_{N}\sim\frac{\kappa_{\infty}}{N},\qquad N\to\infty,$
with a finite positive conductivity $\kappa_{\infty}$.
Our multiscale analysis identifies the mechanism behind this behavior: on each scale $2^{n}$, all path correlations generated by deterministic memory are summarized in a single correction factor $\Delta_{n}$. Once $\Delta_{n}$ is expressed in terms of the increments $\eta_{n}(l+1)-\eta_{n}(l)$, the central feature is the closure hypothesis, based on increasing independence in one sector of the entry/exit phase space and constant strong correlations on the complement. We have computed the effect of the closure of two-point correlation function on the conductivity. This lead to a closed recursion equation for the conductivity where the parameter $\alpha$ gathers the effect of the recollision of trajectories. This parameter is computable numerically with high accuracy.
The limit $\kappa_{\infty}$ is close to the conductivity of the non–backtracking random walk; the small difference quantifies the residual memory not captured by a Markovian model. This suggests applying the same approach to other deterministic systems, such as random Lorentz gases with partial density, deterministic cellular automata with local scatterers, Sinai billiards in finite channels, and transport on random graphs with local deterministic routing rules.
A natural next goal is a rigorous control of the error terms $\eta_{n}(l+1)-\eta_{n}(l)$ under suitable geometric decompositions. Establishing exponential decay in $l$ and $2^{-n}$ would lead to a fully rigorous proof of normal conductivity in the three-dimensional mirrors model.
Data availability
The data and code supporting the findings of this study are available
at https://doi.org/10.5281/zenodo.17578332.
Conflict of interest
The author declares that there is no conflict of interest.
Appendix A Computation of the Single-Slab Crossing Probability $C_{1}$
Let $\gamma=(x_{1},\ldots,x_{n})$ be a path in $\mathbb{Z}^{d}\times{\mathcal{P}}$. The probability of $\gamma$ is
$\mathbb{M}[\gamma]=\left\langle\prod_{i=1}^{n}T(x_{i},x_{i+1};\pi)\right\rangle.$ (25)
Writing $\underline{1}=(1,\ldots,1)$, the crossing probability can be expressed as
$C_{N}=\sum_{\gamma:(\underline{1},\mathbf{e}_{1})\to O_{N}^{+}}\mathbb{M}[\gamma],$ (26)
where the sum is over paths $\gamma=(x_{1},\ldots,x_{n})$ with $x_{1}=(\underline{1},\mathbf{e}_{1})$ and $x_{n}\in O_{N}^{+}$.
For $N=1$ we decompose according to the first step:
$C_{1}$ $=\sum_{\gamma:(\underline{1},\mathbf{e}_{1})\to O_{1}^{+}}\mathbb{M}[\gamma]$
$=\left\langle T((\underline{1},\mathbf{e}_{1}),(\underline{1}+\mathbf{e}_{1},\mathbf{e}_{1});\pi)\right\rangle$
$\quad+\frac{4}{5}\sum_{\gamma:(\underline{1}+\mathbf{e}_{2},\mathbf{e}_{2})\to O_{1}^{+}}\mathbb{M}[\gamma\mid\pi(\underline{1};\mathbf{e}_{1})=\mathbf{e}_{2}]$
$=\frac{1}{5}+\frac{4}{5}\sum_{\gamma:(\underline{1}+\mathbf{e}_{2},\mathbf{e}_{2})\to O_{1}^{+}}\mathbb{M}[\gamma\mid\pi(\underline{1};\mathbf{e}_{1})=\mathbf{e}_{2}].$ (27)
Introducing
$p_{\pm}=\sum_{\gamma:(\underline{1}+\mathbf{e}_{2},\mathbf{e}_{2})\to O_{1}^{\pm}}\mathbb{M}[\gamma\mid\pi(\underline{1};\mathbf{e}_{1})=\mathbf{e}_{2}],$ (28)
we have $p_{+}+p_{-}=1$ because, under the conditioning, there is no loop contained in $\Lambda_{1}$. There is a one-to-one correspondence between paths in $p_{+}$ and $p_{-}$, except for those exiting at $(\underline{1}+\mathbf{e}_{1},\mathbf{e}_{1})$, which are forbidden for $p_{-}$. Thus
$p_{+}=p_{-}+\sum_{\gamma:(\underline{1}+\mathbf{e}_{2},\mathbf{e}_{2})\to(\underline{1}+\mathbf{e}_{1},\mathbf{e}_{1})}\mathbb{M}[\gamma\mid\pi(\underline{1};\mathbf{e}_{1})=\mathbf{e}_{2}],$
and hence
$p_{+}=\frac{1}{2}(1+q),$
with
$q:=\sum_{\gamma:(\underline{1}+\mathbf{e}_{2},\mathbf{e}_{2})\to(\underline{1}+\mathbf{e}_{1},\mathbf{e}_{1})}\mathbb{M}[\gamma\mid\pi(\underline{1};\mathbf{e}_{1})=\mathbf{e}_{2}].$ (29)
We decompose $q=\sum_{n\geq 4}q_{n}$ according to the number of visited vertices.
For $q_{4}$ (paths visiting 4 vertices), there are two such paths, each with conditional probability $(\frac{1}{5})^{3}\frac{1}{3}$, so $q_{4}=(\frac{1}{5})^{3}\frac{2}{3}$. For 6-step loops (visiting 6 vertices) forming rectangles of size $2\times 1$, there are six such loops, each with probability $(\frac{1}{5})^{5}\frac{1}{3}$, giving $q_{6}=6(\frac{1}{5})^{5}\frac{1}{3}$. Thus
$q=0.00597333+\sum_{n\geq 8}q_{n},$
so
$p_{+}\simeq 0.502987,\quad C_{1}\simeq 0.602389.$
Higher $q_{n}$ can be bounded using the analogy with closed trails on the square lattice; each closed orbit of length $n$ has probability at most $(\frac{1}{3})^{n/2}(\frac{1}{5})^{n/2-1}$ and there are at most $3^{n}$ such orbits, so
$q_{n}\leq 5\left(\frac{3}{5}\right)^{\frac{n}{2}}.$
Appendix B Closure hypothesis
Let a volume $\Lambda_{2^{n}}$ and $I^{\pm}$ and $O^{\pm}$ its incoming and outgoing space. We split the incoming and outgoing spaces in $5$ distinct regions on which the correlations functions $\langle t_{n}(x_{1},y_{1})t_{n}(x_{2},y_{2})\rangle$ have a different expressions. We define the time-reversal operation $R(\mathbf{q},\mathbf{p})=(\mathbf{q}-\mathbf{p},-\mathbf{p})$. The $\delta$ factors all come from the bijective and reversible characters of the mirrors model dynamics.
$\Lambda_{2^{n}}$$A^{1}_{uu}$$x_{1}$$y_{1}$$x_{2}$$y_{2}$
$\Lambda_{2^{n}}$$A^{2}_{uu}$$x_{1}$$y_{1}$$x_{2}$$y_{2}$
$\Lambda_{2^{n}}$$A_{cu}$$x_{1}$$y_{1}$$x_{2}$$y_{2}$
$\Lambda_{2^{n}}$$A_{uc}$$x_{2}$$y_{2}$$x_{1}$$y_{1}$
$\Lambda_{2^{n}}$$A_{cc}$$x_{1}$$y_{1}$$x_{2}$$y_{2}$
Figure 7: Schematic boundary configurations and trajectories corresponding to the sets $A^{1}_{uu}$, $A^{2}_{uu}$, $A_{cu}$, $A_{uc}$, and $A_{cc}$. Each panel shows a representative configuration from the defining union, with $x_{i}$ on incoming boundaries $I^{\pm}$ and $y_{i}$ on outgoing boundaries $O^{\pm}$, and two trajectories inside $\Lambda_{2^{n}}$ realizing the corresponding pattern (two u-turns, one crossing and one u-turn, or two crossings).
$A^{1}_{uu}=\{x_{1}\in I^{+},y_{1}\in O^{-},x_{2}\in I^{-},y_{2}\in O^{+}\}\bigcup\{x_{2}\in I^{+},y_{2}\in O^{-},x_{1}\in I^{-},y_{1}\in O^{+}\}$ (30)
$A^{2}_{uu}=\{x_{1}\in I^{+},y_{1}\in O^{-},x_{2}\in I^{+},y_{2}\in O^{-}\}\bigcup\{x_{2}\in I^{+},y_{2}\in O^{-},x_{1}\in I^{+},y_{1}\in O^{-}\}$ (31)
$A_{cu}$ $=$ $\{x_{1}\in I^{+},y_{1}\in O^{+},x_{2}\in I^{-},y_{2}\in O^{+}\}\bigcup\{x_{1}\in I^{+},y_{1}\in O^{+},x_{2}\in I^{+},y_{2}\in O^{-}\}$ (32)
$\bigcup$ $\{x_{1}\in I^{-},y_{1}\in O^{-},x_{2}\in I^{-},y_{2}\in O^{+}\}\bigcup\{x_{1}\in I^{-},y_{1}\in O^{-},x_{2}\in I^{+},y_{2}\in O^{-}\}$
$A_{uc}$ $=$ $\{x_{2}\in I^{+},y_{2}\in O^{+},x_{1}\in I^{-},y_{1}\in O^{+}\}\bigcup\{x_{2}\in I^{+},y_{2}\in O^{+},x_{1}\in I^{+},y_{1}\in O^{-}\}$ (33)
$\bigcup$ $\{x_{2}\in I^{-},y_{2}\in O^{-},x_{1}\in I^{-},y_{1}\in O^{+}\}\bigcup\{x_{2}\in I^{-},y_{2}\in O^{-},x_{1}\in I^{+},y_{1}\in O^{-}\}$
$A_{cc}$ $=$ $\{x_{1}\in I^{+},y_{1}\in O^{+},x_{2}\in I^{+},y_{2}\in O^{+}\}\bigcup\{x_{1}\in I^{+},y_{1}\in O^{+},x_{2}\in I^{-},y_{2}\in O^{-}\}$ (34)
$\bigcup$ $\{x_{1}\in I^{-},y_{1}\in O^{-},x_{2}\in I^{+},y_{2}\in O^{+}\}\bigcup\{x_{1}\in I^{-},y_{1}\in O^{-},x_{2}\in I^{-},y_{2}\in O^{-}\}$
If $(x_{1},y_{1},x_{2},y_{2})\in A^{1}_{uu}$, then
$\langle t_{n}(x_{1},y_{1})t_{n}(x_{2},y_{2})\rangle=c_{n}(x_{1},y_{1})c_{n}(x_{2},y_{2})(1+h^{1}_{n,uu}(x_{1},y_{1},x_{2},y_{2}))$ (35)
If $(x_{1},y_{1},x_{2},y_{2})\in A^{2}_{uu}$, then
$\langle t_{n}(x_{1},y_{1})t_{n}(x_{2},y_{2})\rangle$ $=$ $(1-\delta_{x_{1}x_{2}})(1-\delta_{y_{1}y_{2}})(1-\delta_{x_{1}Ry_{2}})(1-\delta_{x_{2}Ry_{1}})$ (36)
$\cdot$ $c_{n}(x_{1},y_{1})c_{n}(x_{2},y_{2})\left(1+h^{2}_{n,uu}(x_{1},y_{1},x_{2},y_{2})\right)$
$+$ $(\delta_{x_{1}x_{2}}\delta_{y_{1}y_{2}}+\delta_{x_{1}Ry_{2}}\delta_{x_{2}Ry_{1}})c_{n}(x_{1},y_{1})$
If $(x_{1},y_{1},x_{2},y_{2})\in A_{cu}$, then
$\langle t_{n}(x_{1},y_{1})t_{n}(x_{2},y_{2})\rangle=(1-\delta_{Ry_{1}x_{2}})(1-\delta_{y_{1}y_{2}})(1-\delta_{x_{1}x_{2}})(1-\delta_{x_{1}Ry_{2}})$
$\cdot c_{n}(x_{1},y_{1})c_{n}(x_{2},y_{2})\left(1+h_{n,cu}(x_{1},y_{1},x_{2},y_{2})\right)$ (37)
If $(x_{1},y_{1},x_{2},y_{2})\in A_{uc}$, then
$\langle t_{n}(x_{1},y_{1})t_{n}(x_{2},y_{2})\rangle=(1-\delta_{Ry_{1}x_{2}})(1-\delta_{y_{1}y_{2}})(1-\delta_{x_{1}x_{2}})(1-\delta_{x_{1}Ry_{2}})$
$\cdot c_{n}(x_{1},y_{1})c_{n}(x_{2},y_{2})\left(1+h_{n,uc}(x_{1},y_{1},x_{2},y_{2})\right)$ (38)
If $(x_{1},y_{1},x_{2},y_{2})\in A_{cc}$, then
$\langle t_{n}(x_{1},y_{1})t_{n}(x_{2},y_{2})\rangle$ $=$ $(1-\delta_{y_{1}y_{2}})(1-\delta_{x_{1}x_{2}})(1-\delta_{Ry_{2}x_{1}})(1-\delta_{Ry_{1}x_{2}})$ (39)
$\cdot$ $c_{n}(x_{1},y_{1})c_{n}(x_{2},y_{2})\left(1+h_{n,cc}(x_{1},y_{1},x_{2},y_{2})\right)$
$+$ $(\delta_{x_{1}x_{2}}\delta_{y_{1}y_{2}}+\delta_{x_{2}Ry_{1}}\delta_{x_{1}Ry_{2}})c_{n}(x_{1},y_{1})$
All the $h$ function appearing are assumed to have the property that $\sup_{x_{1},y_{1},x_{2},y_{2}}|h_{n,\cdot}(x_{1},y_{1},x_{2},y_{2})|\leq h(n)$, where $h$ is a common positive function such that $h(n)\to 0$ when $n\to\infty$.
The closure assumption used in the main text follows from these structural formulas, together with the observation that the $l=2$ terms in the recursion involve only configurations in $A_{cc}$. The coefficient $\alpha$ is determined by the limiting value of the rescaled diagonal contribution $R_{11}$.
Appendix C Bounds on $R_{11}$, $R_{12}$, and $R_{2}$
We summarize uniform estimates on the terms $R_{11}$, $R_{12}$, and $R_{2}$ defined in the main text.
Lower bound on $R_{11}$. $R_{11}$ can be bounded from below uniformly in $n$ by considering trajectories that, starting at $y_{1}$, remain on the left-most slice of $\Lambda_{2}$ before exiting at $x_{2}\in O_{2}^{-}$.
Fix $y_{1}\in O_{2}^{-}$ and let $q_{\mathrm{L}}$ be the associated leftmost site (so the incoming state into $\Lambda_{2}$ is $(q_{\mathrm{L}},\mathbf{e}_{1})$ by time reversal). We bound $(1-c_{n})^{2}R_{11}$ by summing the squares of the probabilities of disjoint very short return events.
Two-step returns (probability $(2d-1)^{-2}$ each): for $j\in\{2,\dots,d\}$ and $\sigma\in\{\pm 1\}$,
$(q_{\mathrm{L}},\mathbf{e}_{1})\to(q_{\mathrm{L}}+\sigma\mathbf{e}_{j},\sigma\mathbf{e}_{j})\to(q_{\mathrm{L}}+\sigma\mathbf{e}_{j}-\mathbf{e}_{1},-\mathbf{e}_{1}),$
exiting at $x_{2}^{(j,\sigma)}\in O_{2}^{-}$. There are 2(d1)2(d-1) such events, and each gives
$c_{n}\bigl(y_{1},x_{2}^{(j,\sigma)}\bigr)\geq(2d-1)^{-2},$
so
$\sum_{j,\sigma}\bigl(c_{n}(y_{1},x_{2}^{(j,\sigma)})\bigr)^{2}\geq\frac{2(d-1)}{(2d-1)^{4}}.$
Three-step returns (probability $(2d-1)^{-3}$ each): for distinct $j,k\in\{2,\dots,d\}$ and $\sigma,\sigma^{\prime}\in\{\pm 1\}$,
$(q_{\mathrm{L}},\mathbf{e}_{1})\to(q_{\mathrm{L}}+\sigma\mathbf{e}_{j},\sigma\mathbf{e}_{j})\to(q_{\mathrm{L}}+\sigma\mathbf{e}_{j}+\sigma^{\prime}\mathbf{e}_{k},\sigma^{\prime}\mathbf{e}_{k})\to(q_{\mathrm{L}}+\sigma\mathbf{e}_{j}+\sigma^{\prime}\mathbf{e}_{k}-\mathbf{e}_{1},-\mathbf{e}_{1}),$
exiting at $x_{2}^{(j,k,\sigma,\sigma^{\prime})}\in O_{2}^{-}$. There are 4(d1)(d2)4(d-1)(d-2) such events, giving
$\sum_{j\neq k,\sigma,\sigma^{\prime}}\bigl(c_{n}(y_{1},x_{2}^{(j,k,\sigma,\sigma^{\prime})})\bigr)^{2}\geq\frac{4(d-1)(d-2)}{(2d-1)^{6}}.$
Combining we obtain
$R_{11}\geq\frac{2(d-1)}{(2d-1)^{4}}+\frac{4(d-1)(d-2)}{(2d-1)^{6}}.$
In $d=3$,
$R_{11}\geq\frac{4}{5^{4}}+\frac{8}{5^{6}}\approx 6.912\times 10^{-3}.$
Bound on $R_{12}$. $R_{12}$ collects off-diagonal contributions ($y_{2}\neq y_{1}$) in $R_{1}$. Under the closure assumption and using the structural formulas for the two-point function, one shows that each such term is bounded by $O(h(n))$ times a product of $c_{n}$’s, and summing over configurations yields
$R_{12}=O\bigl(h(n)^{2}\bigr)=o(2^{-n}).$
Bound on $R_{2}$. Rewriting (22) and using translation invariance yields
$c_{n}(1-c_{n})^{2}R_{2}$ $=-2\sum_{y_{2}\in O_{1}^{-}}\mathbb{E}[\delta t_{n}(x_{1},y_{1})\delta t_{n}(x_{2},y_{2})]c_{n}(y_{1},x_{2}).$ (40)
Inserting the explicit form of correlations in the appropriate class $A_{cc}$ and discarding configurations where $c_{n}(y_{1},x_{2})=0$ by reversibility, we obtain
$c_{n}(1-c_{n})^{2}|R_{2}|\leq Ch(n)c_{n}^{2},$
for some constant $C>0$, hence
$|R_{2}|\leq\frac{Ch(n)}{2^{n}}.$
Together, these bounds imply the scaling form used in the main text for $\Delta_{n}$ and the resulting asymptotic recursion for $\kappa_{n}$.

### v4

@@
 This leads to strong memory effects and non–Markovian behavior that severely complicate probabilistic analysis.
-In fact, even the question of whether trajectories are typically open or closed is unresolved in $d=3$ at full density.
-Nevertheless, numerical experiments [3] suggest that a remarkable form of universality emerges at large scales: the model exhibits normal conductivity in $d=3$, with a crossing probability close to that of a non–backtracking random walk.
+It is straightforward to see that $\bigl(\tfrac{p}{2d-1}\bigr)^{4}$ provides a lower bound on the probability that a particle starting from the origin becomes trapped in a finite loop forever.
+Proving true diffusivity of the motion in this model is therefore an elusive challenge.
+The best results in this direction are due to Elboim, Gloria, and Hernández [4], who showed that for all dimensions $d\geq 4$ and sufficiently small $p>0$, particle trajectories exhibit diffusive behavior up to times of order $1/p^{\alpha}$, with $\alpha>1$; in particular, trajectories remain open up to this time scale.
+In contrast, in dimension $d=2$ with $p=1$, it is known [6, 2] that every trajectory eventually closes with probability one, while the case $p<1$ was analyzed in [9].
+In the present work we do not address diffusion or recurrence on $\mathbb{Z}^{d}$ itself but instead focus on establishing normal conductivity in the mirrors model.
+This property has been established in a dilute regime for a non-lattice random Lorentz gas [1], and in an anisotropic random lattice Lorentz gas at full density for $d\geq 7$ [7].
+For this purpose, the most fundamental observable is the crossing probability of a finite slab [7, 3].
+Despite the seemingly unfavorable features mentioned above from the viewpoint of diffusion, numerical simulations [3] for the case of full density ($p=1$) in $d=3$ support the conjecture that the mirrors model nevertheless displays normal conductive behavior.
+Specifically, the crossing probability $C_{N}$ of a slab of width $N$ appears to scale as $C_{N}\sim\kappa/N$, with $\kappa\simeq 1.535$, a value remarkably close to $3/2$.
+This coincides with the crossing probability of a non-backtracking random walk, suggesting that despite the strong memory effects inherent to the mirrors dynamics, its large-scale behavior remains close to that of a Markovian process.
 This raises two natural questions: (i) how can a deterministic, non–chaotic system effectively behave as a Markov process at large scales?
@@
 $\langle t_{n}(x_{1},y_{1})t_{n}(x_{2},y_{2})\rangle=\delta_{y_{1}y_{2}}c_{n}(x_{1},y_{1}).$
-The correlation functions represent correlations between two crossings in opposite directions. Using (23), one can show (see Appendix C) that
+Using (23), one can show (see Appendix C) that
 $|R_{2}|\leq\frac{O(h(n))}{2^{n}},$ (24)
@@
 $R_{12}=O\bigl(h(n)^{2}\bigr)=o(2^{-n}).$
-Bound on $R_{2}$. Rewriting (22) and using translation invariance yields
-$c_{n}(1-c_{n})^{2}R_{2}$ $=-2\sum_{y_{2}\in O_{1}^{-}}\mathbb{E}[\delta t_{n}(x_{1},y_{1})\delta t_{n}(x_{2},y_{2})]c_{n}(y_{1},x_{2}).$ (40)
-Inserting the explicit form of correlations in the appropriate class $A_{cc}$ and discarding configurations where $c_{n}(y_{1},x_{2})=0$ by reversibility, we obtain
-$c_{n}(1-c_{n})^{2}|R_{2}|\leq Ch(n)c_{n}^{2},$
-for some constant $C>0$, hence
-$|R_{2}|\leq\frac{Ch(n)}{2^{n}}.$
-Together, these bounds imply the scaling form used in the main text for $\Delta_{n}$ and the resulting asymptotic recursion for $\kappa_{n}$.
+Bound on $R_{2}$.We come now to $R_{2}$.
+By translation invariance, the sum over $x_{3}$ of the last factor produces a factor $c_{n}$ that may be taken out of the sum.
+Next the sum over $y_{2}$ that is running over elements of $O^{+}_{1}$ can be converted into a sum over $y_{2}\in O^{-}_{1}$ with a change of sign in front of the sum.
+This is because thanks to (7) we have the relations
+$\sum_{y_{2}\in O^{+}_{1}}t_{n}(x_{2},y_{2})+\sum_{y_{2}\in O^{-}_{1}}t_{n}(x_{2},y_{2})=1$
+$\sum_{y_{2}\in O^{+}_{1}}c_{n}(x_{2},y_{2})+\sum_{y_{2}\in O^{-}_{1}}c_{n}(x_{2},y_{2})=1$
+The first relation is true for any realizations of the mirrors environment.
+Actually exactly one of the terms is equal to one, all the others are zero.
+The second line follows from the first one by taking expectation.
+So we are left with
+$c_{n}(1-c_{n})^{2}R_{2}$
+$=$ $-2\sum_{y_{2}\in O_{1}^{-}}\mathbb{E}[\delta t_{n}(x_{1},y_{1})\delta t_{n}(x_{2},y_{2})]c_{n}(y_{1},x_{2})$
+Next, let us look at the sum above in more details
+$\sum_{y_{2}\in O_{1}^{-}}\mathbb{E}[\delta t_{n}(x_{1},y_{1})\delta t_{n}(x_{2},y_{2})]c_{n}(y_{1},x_{2})$ (41)
+$=$ $\sum_{y_{2}\in O_{1}^{-}}\left[\mathbb{E}[t_{n}(x_{1},y_{1})t_{n}(x_{2},y_{2})]-c_{n}(x_{1},y_{1})c_{n}(x_{2},y_{2})\right]c_{n}(y_{1},x_{2})$
+Using then the positivity of the $c_{n}$ and the fact that $t_{n}\in\{0,1\}$, we have the easy upper bound $R_{2}\leq\frac{1}{1-c_{n}}$.
+Similarly a lower bound $R_{2}\geq-\frac{c_{n}}{1-c_{n}}$ is obtained so that
+$|R_{2}|\leq\max\left(\frac{1}{1-c_{n}},\frac{c_{n}}{1-c_{n}}\right)$
+Under the closure assumptions detailed above, we now derive a better bound on $R_{2}$.
+Observe first that the sums in (41) involves $(x_{1},y_{1},x_{2},y_{2})\in A_{cc}$.
+Therefore we plug (39) in (41).
+In this expression, we have $c_{n}(y_{1},x_{2})=\langle t_{n}(y_{1},x_{2})\rangle=0$ when $Ry_{1}=x_{2}$ by reversibility and bijectivity of the mirrors map.
+Thus,
+$c_{n}(1-c_{n})^{2}R_{2}=-2\sum_{y_{2}\in O^{-}_{1}}h_{n,cc}(x_{1},y_{1},x_{2},y_{2})(1-\delta_{Ry_{2}x_{1}})c_{n}(x_{1},y_{1})c_{n}(x_{2},y_{2})$
+and
+$c_{n}(1-c_{n})^{2}|R_{2}|$ $\leq$ $4\sum_{y_{2}\in O^{-}_{1}}|h_{n,cc}(x_{1},y_{1},x_{2},y_{2})|c_{n}(x_{1},y_{1})c_{n}(x_{2},y_{2})$
+$\leq$ $O(h(n))c_{n}^{2}$
+and finally
+$|R_{2}|\leq O(h(n))\frac{c_{n}}{(1-c_{n})^{2}}=\frac{O(h(n))}{2^{n}}$